Marco mistakenly made 1,275 defective bikes. The bikes cost $61 each to make. A salvage company decided to purchase the defective bikes as is for $32 each. It would cost Marco $84 per bike to rework the bikes. If Marco rework the bikes, he could sell them for $146 each. No opportunity cost is associated with reworking the bikes. What is the incremental net income from reworking the bikes.
Scrap Rework Sales Rework costs Total
Answer #1
Scrap Rework
Sales (1275*32) = 40,800 (1275*146) = 186,150
Rework cost (1275*84) = 107,100
Total 40,800 79,050
Incremental revenue 38,250

Comment if you face any issues